### v2

@@ 3 Overview of Diffusion Model Technology
Diffusion models (DM) are a class of deep generative models based on probabilistic graphical models. They simulate the random perturbation and deperturbation processes of data [35]to gradually generate high-quality samples from noise. The core idea of diffusion models is to gradually transform data into noise through a forward diffusion process, and then recover the original data from the noise through a reverse diffusion process. This method originated from diffusion processes in non-equilibrium thermodynamics and has achieved breakthrough results in multiple fields such as image generation [19, 36], speech synthesis [37], medical imaging [38], and agricultural remote sensing [39, 40]in recent years. It performs exceptionally well in high-dimensional data processing, image generation, and denoising tasks, particularly demonstrating unique advantages in hyperspectral image processing [40].
@@ 3.1 Basic Principles of Diffusion Models
Diffusion models are a class of deep generative models based on Markov processes, As shown in Figure 3.Their core idea is to simulate the forward diffusion process of data contaminated by noise, then train a neural network to reverse this process and restore the original data, i.e., the reverse process, thereby achieving controllable generation from noise to data.
@@ 3.1.1 Forward Diffusion Process
Given the initial sample $\mathrm{x}_{0}\thicksim q(\mathrm{x}_{0})$, the forward diffusion process is defined as gradually adding Gaussian noise over T time steps, in the form:
@@
Where $\beta_{t}\in(0,1)$ is the noise variance scheduling parameter for time step t, which is typically applied linearly, cosine, or learning-based across the entire time series.
This process constitutes a Markov chain. After t cumulative executions, the intermediate state $x_{t}$ at step $t$ can be directly sampled from $x_{0}$, satisfying:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{0})=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathrm{x},(1-\bar{\alpha}_{t})I\right)$ (2)
Where,The parameter $\alpha_{t}$ represents the cumulative factor for retaining image information and is defined as:
@@
Where,The above formula shows that during training, it is possible to skip explicit sampling at each step of the Markov chain and instead generate $x_{t}$ at any time step from $x_{0}$, greatly simplifying the training process.
To intuitively understand the evolution path of the diffusion process, the following diagram can be used: Each step of noise addition is slight, but as time t increases, the image structure and semantic information are gradually lost, eventually approaching an isotropic standard Gaussian distribution.
@@ 3.2 Denoising Diffusion Model (DDPM)
$E_{q}\left[-\log p_{\theta}(x_{0})\right]\leq E_{q}\left[D_{KL}\left(q(x_{T}|x_{0})\|p(x_{T})\right)\right]+\sum_{t>1}D_{KL}\left(q(x_{t-1}|x_{t},x_{0})\|p_{\theta}(x_{t-1}|x_{t})\right)-\log p_{\theta}(x_{0}|x_{1})$ (8)
@@
$\mathcal{L}_{\mathrm{simple}}=$ (11)
$\mathbb{E}_{\mathcal{X}_{0},\boldsymbol{\epsilon}}\left[\frac{\beta_{t}^{2}}{2\sigma_{\mathrm{T}}^{2}\alpha_{\mathrm{t}}(1-\bar{\alpha}_{\mathrm{t}})}\right]\parallel\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{\mathrm{t}}}x_{0}+\sqrt{1-\bar{\alpha}_{\mathrm{t}}}\epsilon)\parallel^{2}$
@@ 4.7 Small Sample Data Image Generation
Figure 12: Shows seven areas of application for diffusion models in agriculture: (1) data augmentation, (2) image generation, (3) remote sensing and hyperspectral image reconstruction, (4) pest and disease diagnosis and decision support, (5) image generation from small sample data, (6) multimodal fusion and agricultural knowledge generation, and (7) agricultural image classification. No abbreviation, first and last two initials
@@ 5.2.3 Pest and disease diagnosis and auxiliary decision-making experiments
This experiment evaluated the performance of different models in sunflower disease detection using metrics such as precision, recall, accuracy, F1 score, and mean average precision (mAP@75). The results showed that the models achieved significant improvements through architectural optimization and task adaptation, particularly in complex background and limited sample scenarios. Table 6 shows that RT-DETRv2 performed the worst, as it focuses on global features and lacks the ability to extract fine details; Mask R-CNN achieved good results through instance segmentation, but still has room for improvement in small spot segmentation; the DETR model performed well, but faced challenges in background noise and small spot detection; The LeafDetection[130] model, specifically designed for sunflower disease detection, demonstrates a clear advantage in fine-grained feature extraction, outperforming other models in all metrics with an accuracy of 0.94, recall of 0.92, precision of 0.93, F1 score of 0.93, and mAP@75 is 0.92, proving the effective combination of deep learning technology and customized disease feature modeling. According to the article, the method proposed by the authors can be named "Few-Shot Diffusion Detection Network (FSDDN)". This name highlights the application of few-shot learning and diffusion models for high-quality disease detection in small sample situations.
@@ 5.2.4 Multimodal fusion experiment
Model Precision Recall Accuracy mAP F1 Score FPS Params (M) FLOPs (G) Memory (GB)
YOLOv9 0.83 0.79 0.81 0.82 0.81 35 57.3 189 10
Retinanet 0.85 0.81 0.83 0.84 0.83 25 52 198 8
Efficientdet 0.87 0.83 0.85 0.86 0.85 30 77 410 12
YOLOv10 0.89 0.85 0.86 0.87 0.87 40 29.5 160 9
DETR 0.91 0.87 0.88 0.89 0.89 22 60 253 11
JuDiformer 0.93 0.89 0.90 0.91 0.91 28 61.32 207 10

### v3

@@ 3 Overview of Diffusion Model Technology
Diffusion models (DMs) are a class of deep generative models based on probabilistic graphical models. They simulate the processes of random perturbation and denoising of data [35], gradually generating high-quality samples from pure noise. The core concept involves transforming data into noise through a forward diffusion process and recovering the original data via a learned reverse process. Originating from diffusion principles in non-equilibrium thermodynamics, DMs have achieved remarkable success in fields such as image generation [19, 36], speech synthesis [37], medical imaging [38], and agricultural remote sensing [39, 40]. They excel in processing high-dimensional data, particularly in image synthesis and denoising tasks, and have shown distinct advantages in hyperspectral image analysis [40].
@@ 3.1 Basic Principles of Diffusion Models
Diffusion models rely on Markov processes to simulate data transformation. As illustrated in Figure 4, the model defines a forward process where data is progressively corrupted with Gaussian noise and a reverse process that restores the data using a neural network trained to invert the noise.
@@ 3.1.1 Forward Diffusion Process
Given an initial sample $\mathrm{x}_{0}\sim q(\mathrm{x}_{0})$, the forward diffusion process introduces Gaussian noise over $T$ time steps, defined as:
@@
Here, $\beta_{t}\in(0,1)$ is the noise variance schedule at timestep $t$, which may follow linear, cosine, or learnable schemes.
This forms a Markov chain. After $t$ iterations, the intermediate state $x_{t}$ can be directly sampled from $x_{0}$ as:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{0})=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{\bar{\alpha}_{t}}\,\mathrm{x}_{0},(1-\bar{\alpha}_{t})I\right)$ (2)
Here, $\bar{\alpha}_{t}$ is the cumulative product of retained information up to time $t$:
@@
This formulation enables direct sampling of $x_{t}$ from $x_{0}$ at any timestep, avoiding explicit sequential sampling and simplifying training.
To visualize the diffusion process, consider that each noise addition introduces minor perturbations. As $t$ increases, structural and semantic content in the image degrades, eventually resulting in an isotropic standard Gaussian distribution.
@@ 3.2 Denoising Diffusion Model (DDPM)
$E_{q}\left[-\log p_{\theta}(x_{0})\right]$ $\leq E_{q}\left[D_{KL}\left(q(x_{T}|x_{0})\|p(x_{T})\right)\right]$ (8)
$+\sum_{t>1}D_{KL}\left(q(x_{t-1}|x_{t},x_{0})\|p_{\theta}(x_{t-1}|x_{t})\right)-\log p_{\theta}(x_{0}|x_{1})$
@@
$\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathcal{X}_{0},\boldsymbol{\epsilon}}\left[\frac{\beta_{t}^{2}}{2\sigma_{\mathrm{T}}^{2}\alpha_{\mathrm{t}}(1-\bar{\alpha}_{\mathrm{t}})}\right]\parallel\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{\mathrm{t}}}x_{0}+\sqrt{1-\bar{\alpha}_{\mathrm{t}}}\epsilon)\parallel^{2}$ (11)
@@ 4.7 Small Sample Data Image Generation
Figure 13: Shows seven areas of application for diffusion models in agriculture: (1) data augmentation, (2) image generation, (3) remote sensing and hyperspectral image reconstruction, (4) pest and disease diagnosis and decision support, (5) image generation from small sample data, (6) multimodal fusion and agricultural knowledge generation, and (7) agricultural image classification. No abbreviation, first and last two initials.
@@ 5.2.3 Pest and disease diagnosis and auxiliary decision-making experiments
This experiment evaluated the performance of different models in sunflower disease detection using metrics such as precision, recall, accuracy, F1 score, and mean average precision (mAP@75). The results showed that the models achieved significant improvements through architectural optimization and task adaptation, particularly in complex background and limited sample scenarios. Table 6 shows that RT-DETRv2 performed the worst, as it focuses on global features and lacks the ability to extract fine details; Mask R-CNN achieved good results through instance segmentation, but still has room for improvement in small spot segmentation; the DETR model performed well, but faced challenges in background noise and small spot detection; The LeafDetection[130] model, specifically designed for sunflower disease detection, demonstrates a clear advantage in fine-grained feature extraction, outperforming other models in all metrics with an accuracy of 0.94, recall of 0.92, precision of 0.93, F1 score of 0.93, and mAP@75 is 0.92, proving the effective combination of deep learning technology and customized disease feature modeling. According to the article, the method proposed by the authors can be named ”Few-Shot Diffusion Detection Network (FSDDN)”. This name highlights the application of few-shot learning and diffusion models for high-quality disease detection in small sample situations.
@@ 5.2.4 Multimodal fusion experiment
Metric YOLOv9 Retinanet Efficientdet YOLOv10 DETR JuDiformer
Precision 0.83 0.85 0.87 0.89 0.91 0.93
Recall 0.79 0.81 0.83 0.85 0.87 0.89
Accuracy 0.81 0.83 0.85 0.86 0.88 0.90
mAP 0.82 0.84 0.86 0.87 0.89 0.91
F1 Score 0.81 0.83 0.85 0.87 0.89 0.91
FPS 35 25 30 40 22 28
Params (M) 57.3 52 77 29.5 60 61.32
FLOPs (G) 189 198 410 160 253 207
Memory (GB) 10 8 12 9 11 10